# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### v2

@@ Abstract
We present EE-LLM, a framework for large-scale training and inference of early-exit large language models (LLMs). While recent works have shown preliminary evidence for the efficacy of early exiting in accelerating LLM inference, EE-LLM makes a foundational step towards scaling up early-exit LLMs by supporting their training and inference with massive 3D parallelism. Built upon Megatron-LM, EE-LLM implements a variety of algorithmic innovations and performance optimizations tailored to early exiting, including a lightweight method that facilitates backpropagation for the early-exit training objective with pipeline parallelism, techniques of leveraging idle resources in the original pipeline schedule for computation related to early-exit layers, and two approaches of early-exit inference that are compatible with KV caching for autoregressive generation. Our analytical and empirical study shows that EE-LLM achieves great training efficiency with negligible computational overhead compared to standard LLM training, as well as outstanding inference speedup without compromising output quality. To facilitate further research and adoption, we release EE-LLM at https://github.com/pan-x-c/EE-LLM. [^1] We will continuously update the codebase and arXiv version.
11footnotetext: Co-first authors. 22footnotetext: Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have amazed the world with their astonishing abilities and performance in solving a wide range of problems [8, 50, 11, 89, 74, 75]. This is accompanied by excessive costs and carbon emissions for training and deploying these models, as their sizes grow rapidly in recent years. In general, costs for inference are dominant in the long run, as each model will be deployed to solve many problems for a long period of time. This has inspired researchers and engineers to develop various approaches for accelerating LLM inference.
@@
The focus of this work is early exiting, which accelerates inference by allowing a deep neural network to make predictions and exit early in the network for certain inputs, without running the forward pass through the full network. This is achieved by augmenting a standard neural network architecture (with a single exit at the end) with additional early-exit layers that transform intermediate hidden states into early outputs. The early-exit model architecture, as visualized in Figure 1, not only retains the full capacity of a large model, but is also capable of adaptively using a smaller amount of computation for solving simpler problems. The idea of early exiting is a natural analogue of how human speaks, thinks, and make decisions: not every problem requires or deserves the same amount of consideration, and one shall opt for fast reaction to simple problems without overthinking [31]. Early exiting has been an active research area and widely applied in natural language processing, computer vision, and other areas [23, 45, 25, 94, 19, 63, 81, 62, 82, 42, 26, 73, 27, 31, 38, 60, 24, 83, 14]. More recently, it starts to gain attention in the LLM domain [62, 13, 3, 76], and is recognized as a promising direction for further reducing the latency and costs of LLM inference [53].
1.1 Goal and motivations
@@
The primary goal of this work is to build the infrastructure for scaling up training and inference of early-exit LLMs. This is motivated by the observation that the sizes of early-exit models in prior works are still relatively small. The largest early-exit LLM that we are aware of is a Llama 2 model [75] with 13 billion (13B) parameters [76]; in contrast, standard LLMs at much larger scales, e.g. the 175B GPT-3 [8], 530B Megatron-Turing NLG [68], 540B PaLM [11], or even larger sparsely activated models, have been well trained and deployed in many applications. It is an urgent need for the community to truly understand the efficacy of early exiting for LLMs at larger scales, which is indispensable for making early exiting a useful and practical option in complex scenarios that only sufficiently large LLMs can handle.
1.2 Challenges
@@
The first and foremost question is how to train an early-exit LLM that is too large to fit into the memory of one single device (e.g. GPU). While state-of-the-art frameworks like Megatron-LM [67, 49, 36], DeepSpeed [58, 68], Mesh-TensorFlow [65], Alpa [92], InternLM [72], and many more, support training standard LLMs at large scales with data parallelism and model parallelism (including tensor, sequence and pipeline parallelism), they do not provide native support for early-exit LLMs. One particular challenge lies in pipeline parallelism [47, 48, 20, 41], which partitions the model along the depth dimension into multiple pipeline stages, connected by limited point-to-point communication between devices; this seems to contradict with early-exit models, as the early-exit training loss is typically an aggregation of losses for multiple (early or final) exits that are now located separately on different pipeline stages. Despite the necessity of pipeline parallelism in many scenarios, we are not aware of any implementation that supports training early-exit LLMs with pipeline parallelism.
@@
Moreover, training efficiency for early-exit LLMs requires special design. While sizes of early-exit layers are often regarded as negligible for many neural network architectures, this is not the case for LLMs, where each early exit contains (at least) a large output embedding matrix that transforms hidden states into logits on the vocabulary. A naive implementation of early-exit LLM training can cause large computational overhead compared to standard LLM training.
@@
Finally, with regards to autoregressive generation tasks (where tokens are generated one by one, depending on previously generated tokens via the attention mechanism), a naive implementation of early-exit inference is not compatible with KV caching, a standard technique of storing the keys and values of previously generated tokens at each layer. More specifically, if the current token is generated via early exiting at some layer, then its KV caches in later layers are missing, which hinders the generation of future tokens. Given that KV caching is enabled by default in most cases, the efficacy of early exiting for autoregressive generation might be questionable if its conflict with KV caching is not well resolved.
1.3 Main contributions
@@
We propose EE-LLM, a system for large-scale training and inference of early-exit (EE) LLMs with 3D parallelism, which is designed to tackle the aforementioned challenges. EE-LLM is built upon Megatron-LM [67, 49, 36, 68], and augments it with various functionalities for early exiting. In addition to compatibility with existing functionalities of 3D parallelism provided by Megatron-LM, EE-LLM also implements a variety of algorithmic innovations, including a lightweight method that facilitates backpropagation for the early-exit training objective through pipeline stages, various techniques of leveraging idle resources in the original pipeline schedule for computation related to early-exit layers, and two approaches of early-exit inference that are compatible with KV caching (with one based on a novel form of pipeline parallelism and another based on KV recomputation). Implementation of EE-LLM has been well optimized for maximum training and inference efficiency. Our analytical and empirical study confirms that, with negligible computational overhead (caused by early-exit layers) during training with 3D parallelism, one obtains an early-exit LLM that generates tokens with adaptive token-wise exit selection, achieving outstanding speedup without compromising output quality during inference. In other words, EE-LLM facilitates training and inference of early-exit LLMs that are as large as the maximum sizes of standard LLMs allowed by Megatron-LM, given the same amount of computational resources. The source code for EE-LLM can be found at https://github.com/pan-x-c/EE-LLM.
@@
Figure 1: The model architecture of an early-exit LLM. New components related to early exiting, which are absent from a standard LLM, are highlighted in blue color. In this figure, each $\bm{\theta}_{i}$ represents a sequence of Transformer layers in the backbone of the LLM (with some additional modules in $\bm{\theta}_{1}$ for input processing); in addition, each $\bm{\phi}_{i}$ represents an early-exit or final-output layer that converts the hidden state $\bm{x}_{i}$ into an output $\bm{o}_{i}$, e.g. logits for next-token prediction.
2 Preliminaries
2.1 Transformers
The Transformer architecture [77, 71] has been playing a dominant role in natural language processing (NLP) and large language models (LLMs) [7, 90]. It is typically composed of an input embedding layer, a stack of Transformer layers, and finally an output layer. Each Transformer layer consists of cross-attention and/or self-attention modules [4, 34, 52], a multi-layer perceptron (MLP), and layer normalization (LayerNorm [2] or RMSNorm [88]). Transformers can be categorized into three types: encoder-only, encoder-decoder, and decoder-only. For the latter two, there is an output embedding matrix in the output layer, which transforms hidden states into logits on a (typically large) vocabulary that can be used for generating tokens. An LLM can be learned by unsupervised pre-training, e.g. minimizing the negative log-likelihood of next-token prediction on a large corpus [55, 56]. In this work, we focus on the decoder-only generative pre-training (GPT) Transformer architecture [55, 56], though many of our ideas are widely applicable to other Transformer architectures or generic deep neural networks.
2.2 Early-exit LLMs
As mentioned earlier, an early-exit neural network can be obtained by adding to a standard nerual network some early-exit layers that turn intermediate hidden states into early outputs [81, 63]. During inference for a given input, the model starts a forward pass and decides (at each early exit) whether to return an output or continue forwarding via certain rules, e.g. to return an output whenever the confidence of prediction is above a pre-defined threshold [63, 61].
The standard way of training an early-exit model is to minimize a weighted sum of early-exit and final-exit training losses [63, 61]. Note that early-exit layers bring additional computational overhead to training. This is especially the case for LLMs, primarily due to the large output embedding matrix of size $h\times V$ within each early-exit layer, where $h$ is the hidden dimension and $V$ is the vocabulary size. We call an early-exit layer minimalistic if it has the same structure as the final output layer of the GPT model architecture [55, 56], which includes an output embedding matrix, plus an optional LayerNorm/RMSNorm in front of it. Additional modules can be added to early-exit layers for increased expressivity and adaptivity of early exits.
2.3 Megatron-LM and 3D parallelism
Megatron-LM [67, 49, 36, 68] is one of the state-of-the-art systems for large-scale training of LLMs with 3D parallelism on a massive number of GPUs. With data parallelism, each GPU handles the forward and backward computation for one part of the data batch, and then the results are aggregated at the end of the training iteration. When the model is too large to fit in a single GPU, model partitioning becomes necessary and can be used in conjunction with data parallelism. With tensor (and sequence) parallelism, each large module (e.g. a linear layer) is divided into multiple pieces that are assigned to different GPUs, so that each computational task related to it (e.g. large matrix multiplication) can be divided into smaller tasks and solved in parallel. One major limitation with tensor (and sequence) parallelism is that it requires expensive collective communication such as all-reduce operations, and thus is only viable for high-end GPUs within the same computing node, with high-bandwidth communication among them.
Pipeline parallelism [47, 48, 20, 41], on the other hand, partitions a deep model along the depth dimension into multiple pipeline stages. Moreover, each data batch is divided into multiple microbatches, and their forward/backward computational tasks are scheduled among those multiple pipeline stages. More specifically, each stage performs the forward computation for each microbatch and sends the resulted hidden states to another stage; later on, it performs the backward computation for the same microbatch after receiving the gradients of the training objective with respect to the sent hidden states. Pipeline parallelism only requires sparse and inexpensive point-to-point (P2P) communication between pipeline stages, which makes it applicable and oftentimes must-have in much broader scenarios when tensor (and sequence) parallelism is infeasible or insufficient, whether in GPU clusters or in decentralized settings [86, 84]. The main concern with pipeline parallelism is its low utilization rate of computational resources, due to pipeline bubbles and load imbalance across pipeline stages [49]; this will be discussed in more details in Section 4.2.
@@ 3 An overview of EE-LLM
This section provides an overview of our system for scaling up sizes, training and inference of early-exit LLMs, with flexible configurations and a wide range of functionalities. More details about training and inference will be introduced in Sections 4 and 5, respectively. Implementation details can be found in Section 6.
3.1 Model architectures
@@
We implement in EE-LLM an early-exit Transformer architecture (which is built upon the GPT Transformer architecture [55, 56] originally implemented in Megatron-LM), with support for various configurations. In particular, users can (1) specify arbitrary layers to add early exits to; (2) add trainable modules to each early-exit layer (between the hidden states on the backbone and the minimalistic early-exit output head), such as an MLP or a complete Transformer layer; and (3) choose to tie [54, 61, 76] or untie the input and output embedding matrices of all early-exit/final-exit layers. Each particular configuration has its own pros and cons, as will be discussed in later sections. With this in mind, EE-LLM has been designed to cover a wide range of common configurations, so that users can easily try them out and choose the most suitable ones for their own use cases.
3.2 Training
@@
EE-LLM contains the essential functionalities for training early-exit LLMs, which tackle the main challenges outlined in Section 1.2, i.e. how to train with 3D parallelism (especially pipeline parallelism) while minimizing the computational overhead (compared to standard LLM training) caused by early-exit layers. In addition to substantial engineering efforts for compatibility with existing functionalities in Megatron-LM, we design and implement a simple yet effective algorithm that facilitates pipeline parallelism with multiple early-exit or final-output training losses located on different pipeline stages, which is not possible in standard pipeline parallelism implemented by Megatron-LM or other frameworks. Moreover, our analytical and empirical study shows that training an early-exit LLM with EE-LLM is almost as efficient as training a standard LLM, in terms of training time and peak GPU memory; this is achieved by various performance optimizations that we design and implement in EE-LLM, especially the ones that passively or actively leverage idle computational resources in pipeline parallelism. Finally, EE-LLM contains some advanced features for more fine-grained control or optimization of the training procedure, including the option of changing early-exit loss weights during training, and a novel method of further improving resource utilization by filling pipeline bubbles with useful computation.
3.3 Inference
@@
We design and implement two methods to tackle the major challenge of early-exit LLM inference for autoregressive generation, namely the conflict between early exiting and KV caching (as explained in Section 1.2). One method is based on KV recomputation, which runs the forward pass with a list of recent tokens when generating each token, and can be regarded as a variant of synchronized parallel decoding recently proposed in [3]. The other method is based on a novel form of pipeline parallelism, which parallelizes the forward pass of the current token at a certain pipeline stage with some KV-related computation (if any) of previous tokens at later stages.
@@ 4 Training
In this section, we first introduce (in Section 4.1) the essentials of scaling up early-exit LLM training with massive 3D parallelism. In particular, we demonstrate how to leverage pipeline parallelism via a novel approach of executing backpropagation for the early-exit training loss through pipeline stages. Next, we show (in Section 4.2) how to maximize training efficiency via various performance optimizations, e.g. choosing wisely where to add early exits, and scheduling computation in the correct order. By leveraging idle computational resources in pipeline parallelism, training an early-exit LLM bears negligible overhead compared to training a standard LLM, as confirmed by our analytical and empirical study. Finally, we explore (in Section 4.3) some advanced features implemented in EE-LLM, including the option to change the weights of early-exit losses during the training process, and a novel approach of improving hardware utilization by filling pipeline bubbles with useful computation for additional microbatches.
@@ 4.1 Backpropagation through pipeline stages
Our implementation of training, i.e. optimizing the loss function in Eq. (1), is compatible with all types of parallelism in Megatron-LM. Indeed, with some engineering efforts, existing functionalities in Megatron-LM for data and tensor/sequence parallelism are directly applicable. The major challenge lies in pipeline parallelism, since it is not immediately clear how to calculate gradients for Eq. (1) via backpropagation through pipeline stages. In a single-GPU scenario with vanilla PyTorch, one simply needs to define loss as the weighted sum of all early-exit and final-exit losses, and then run loss.backward() for gradient calculation. This is obviously not feasible with pipeline parallelism, since losses $\{{\mathcal{L}}_{i}^{\mathsf{exit}}\}$ are located on different pipeline stages, and there is only limited P2P communication between each pair of adjacent stages (as explained in Section 2.3). On the other hand, Megatron-LM only supports backpropagation for a single loss function defined in the last pipeline stage.

### v3

@@ Abstract
We present EE-LLM, a framework for large-scale training and inference of early-exit large language models (LLMs). While recent works have shown preliminary evidence for the efficacy of early exiting in accelerating LLM inference, EE-LLM makes a foundational step towards scaling up early-exit LLMs by supporting their training and inference with massive 3D parallelism. Built upon Megatron-LM, EE-LLM implements a variety of algorithmic innovations and performance optimizations tailored to early exiting, including a lightweight method that facilitates backpropagation for the early-exit training objective with pipeline parallelism, techniques of leveraging idle resources in the original pipeline schedule for computation related to early-exit layers, and two approaches of early-exit inference that are compatible with KV caching for autoregressive generation. Our analytical and empirical study shows that EE-LLM achieves great training efficiency with negligible computational overhead compared to standard LLM training, as well as outstanding inference speedup without compromising output quality. To facilitate further research and adoption, we release EE-LLM at https://github.com/pan-x-c/EE-LLM.
Large Language Models, Early Exiting, 3D Parallelism
@@ 1 Introduction
Large language models (LLMs) have amazed the world with their astonishing abilities and performance in solving a wide range of problems (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2023; Zhang et al., 2022; Touvron et al., 2023a, b). This is accompanied by excessive costs and carbon emissions for training and deploying these models, as their sizes have grown rapidly in recent years. In general, costs for inference are dominant in the long run, as each model will be deployed to solve many problems for a long period of time. This has inspired researchers and engineers to develop various approaches for accelerating LLM inference.
@@
The focus of this work is early exiting, which accelerates inference by allowing a deep neural network to make predictions and exit early in the network for certain inputs, without running the forward pass through the full network. This is achieved by augmenting a standard neural network architecture (with a single exit at the end) with additional early-exit layers that transform intermediate hidden states into early outputs. The early-exit model architecture, as visualized in Figure 1, not only retains the full capacity of a large model, but is also capable of adaptively using a smaller amount of computation for solving simpler problems. The idea of early exiting is a natural analogue of how humans speak, think, and make decisions: not every problem requires or deserves the same amount of consideration, and one shall opt for fast reaction to simple problems without overthinking (Kaya et al., 2019). Early exiting has been an active research area and widely applied in natural language processing, computer vision, and other areas (Liu et al., 2020; Elbayad et al., 2020; Schwartz et al., 2020; Xin et al., 2020; Schuster et al., 2021; Teerapittayanon et al., 2016; Huang et al., 2018; Scardapane et al., 2020; Han et al., 2022; Xu & McAuley, 2023). More recently, it has started to gain attention in the generative LLM domain (Schuster et al., 2021; Corro et al., 2023; Bae et al., 2023; Varshney et al., 2023), and is recognized as a promising direction for further reducing the latency and costs of LLM inference (Pope et al., 2022).
Goal and motivations.
@@
Our primary goal is to build the infrastructure for scaling up training and inference of early-exit LLMs. This is motivated by the observation that sizes of early-exit models in prior works are still relatively small. While the largest early-exit LLM that we are aware of has 13 billion parameters (Varshney et al., 2023), standard LLMs at much larger scales, e.g. the 175B GPT-3 (Brown et al., 2020), 530B Megatron-Turing NLG (Smith et al., 2022), 540B PaLM (Chowdhery et al., 2023), or even larger sparsely activated models, have been well trained and deployed in many applications. It is an urgent need for the community to truly understand the efficacy of early exiting for LLMs at larger scales, which is indispensable for making early exiting a useful and practical option in complex scenarios that only sufficiently large LLMs can handle.
Challenges.
@@
The first and foremost question is how to train an early-exit LLM that is too large to fit into the memory of one single device (e.g. GPU). While state-of-the-art frameworks like Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b), DeepSpeed (Rasley et al., 2020; Smith et al., 2022), Alpa (Zheng et al., 2022), and many more, support training standard LLMs at large scales with data parallelism and model parallelism (including tensor, sequence, and pipeline parallelism), they do not provide native support for early-exit LLMs. One particular challenge lies in pipeline parallelism (Narayanan et al., 2019, 2021a; Fan et al., 2021), which partitions the model along the depth dimension into multiple pipeline stages, connected by limited point-to-point communication between devices; this seems to contradict training early-exit models, as the training objective is typically an aggregation of losses for multiple exits that are now located separately on different pipeline stages. Despite the necessity of pipeline parallelism in many scenarios, we are not aware of any implementation that supports training early-exit LLMs with it.
@@
Moreover, training efficiency for early-exit generative LLMs requires special design, since each early exit contains (at least) a large output embedding matrix that transforms hidden states into logits on the vocabulary, which can constitute a non-trivial proportion of the whole model. Naive implementation of early-exit LLM training can cause large computational overhead compared to standard LLM training.
@@
Finally, with regard to autoregressive generation (where tokens are generated one by one, depending on previously generated tokens via the attention mechanism), naive implementation of early-exit inference is not compatible with KV caching, a standard technique of storing the keys and values of previously generated tokens at each layer. Indeed, if the current token is generated via early exiting at some layer, then its KV caches in later layers are missing, which hinders the generation of future tokens. Given that KV caching is enabled by default in most cases, the efficacy of early exiting for autoregressive generation might be questionable if its conflict with KV caching is not well resolved.
Main contributions.
@@
We propose EE-LLM, a system for large-scale training and inference of early-exit (EE) LLMs with 3D parallelism, which is designed to tackle the aforementioned challenges. EE-LLM is built upon Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; Smith et al., 2022), and augments it with various functionalities for early exiting. In addition to compatibility with existing functionalities of 3D parallelism provided by Megatron-LM, EE-LLM also implements a variety of algorithmic innovations, including a lightweight method that facilitates backpropagation for the early-exit training objective through pipeline stages, various techniques of leveraging idle resources in the original pipeline schedule for computation related to early-exit layers, and two approaches of early-exit inference that are compatible with KV caching. Implementation of EE-LLM has been well optimized for maximum training and inference efficiency. Our analytical and empirical study confirms that, with negligible computational overhead caused by early-exit layers during training with 3D parallelism, one obtains an early-exit LLM that generates tokens with adaptive token-wise exit selection, achieving outstanding inference speedup without compromising output quality. With EE-LLM, it is now possible to train and deploy early-exit LLMs that are as large as the maximum sizes of standard LLMs allowed by Megatron-LM, given the same amount of computational resources. The source code for EE-LLM can be found at https://github.com/pan-x-c/EE-LLM.
@@
Figure 1: The model architecture of an early-exit LLM. Additional components compared to a standard LLM are highlighted in blue. Each $\bm{\theta}_{i}$ represents a sequence of Transformer layers in the backbone of the LLM, with some additional modules in $\bm{\theta}_{1}$ for input processing. Each $\bm{\phi}_{i}$ represents an early or final-exit layer that converts hidden states $\bm{x}_{i}$ into output $\bm{o}_{i}$, e.g. logits for next-token prediction.
Organization.
Section 2 provides a high-level overview of EE-LLM, while Sections 3 and 4 focus on training and inference, respectively. The efficacy of EE-LLM is validated by numerical experiments in Section 5. The appendix includes extended preliminaries and related works, additional experiments, as well as more details about the training efficiency, advanced features, and implementation of EE-LLM.
@@ 2 An overview of EE-LLM
This section provides an overview of our system for scaling up sizes, training and inference of early-exit LLMs, with flexible configurations and a wide range of functionalities.
Model architectures.
@@
We implement in EE-LLM an early-exit Transformer architecture, which is built upon the generative pre-training (GPT) Transformer architecture (Radford et al., 2018, 2019) originally implemented in Megatron-LM. EE-LLM allows users to (1) specify arbitrary layers to add early exits to; (2) add trainable modules to each early-exit layer, e.g. a multi-layer perceptron (MLP) or a complete Transformer layer, on top of the minimalistic structure (with an output embedding matrix, plus an optional layer normalization module in front of it); and (3) choose to tie (Press & Wolf, 2017; Schuster et al., 2022; Varshney et al., 2023) or untie the input and output embedding matrices of all early/final-exit layers. Each option has its own pros and cons, as will be discussed in later sections. With this in mind, EE-LLM has been designed to cover a wide range of configurations, so that users can easily try them out and choose the most suitable ones for their own use cases.
Training.
@@
EE-LLM contains the essential functionalities for training early-exit LLMs, which tackle the main challenges outlined in Section 1, i.e. how to train with 3D parallelism while minimizing the computational overhead caused by early-exit layers. In addition to substantial engineering efforts for compatibility with existing functionalities in Megatron-LM, we design and implement a simple yet effective algorithm that facilitates pipeline parallelism with multiple early/final-exit training losses located on different pipeline stages, which is not possible in standard pipeline parallelism implemented by Megatron-LM or other frameworks. Moreover, our analytical and empirical study shows that training an early-exit LLM with EE-LLM is almost as efficient as training a standard LLM, in terms of training time and peak GPU memory. This is achieved by various performance optimizations that we design and implement in EE-LLM, especially the ones that leverage idle computational resources in standard pipeline parallelism. Finally, EE-LLM contains some advanced features for fine-grained control or optimization of the training process, including the option of changing early-exit loss weights during training, and a novel method of further improving resource utilization by filling pipeline bubbles with useful computation.
Inference.
@@
We design and implement two methods to tackle the major challenge of early-exit LLM inference for autoregressive generation, namely the conflict between early exiting and KV caching (as explained in Section 1). One method is based on KV recomputation, which runs the forward pass with a batch of recent tokens when generating each token. The other method is based on a novel form of pipeline parallelism, which parallelizes the forward pass of the current token at a certain pipeline stage with some KV-related computation of previous tokens at later stages.
@@ 3 Training
We first present in Section 3.1 the essentials of scaling up early-exit LLM training with 3D parallelism, including a novel approach to executing backpropagation for the early-exit training objective through pipeline stages. Then, we analyze the training efficiency achieved by EE-LLM in Section 3.2, and explore some advanced features in Section 3.3.
@@ 3.1 Backpropagation through pipeline stages
Our implementation for optimizing the loss function in Eq. (1) is compatible with all types of parallelism in Megatron-LM. Indeed, with some engineering efforts, existing functionalities in Megatron-LM for data and tensor/sequence parallelism are directly applicable. The major challenge lies in pipeline parallelism, since it is not immediately clear how to calculate gradients for Eq. (1) via backpropagation through pipeline stages. In a single-GPU scenario with vanilla PyTorch, one simply needs to define loss as the weighted sum of losses at all exits, and then run loss.backward() for gradient calculation. This is not feasible with pipeline parallelism, since losses $\{{\mathcal{L}}_{i}^{\mathsf{exit}}\}$ are now located on different GPUs, and there is only limited P2P communication between each pair of adjacent stages. On the other hand, Megatron-LM only supports backpropagation for a single loss function defined in the last stage.